# Possible Unconventional Surface Superconductivity in the Half-Heusler YPtBi

### v1

@@ Acknowledgements.
We thank Ronny Thomale and Jörg Schmalian for illuminating discussions. This work was supported by the U. S. Department of Energy (DOE) Office of Basic Energy Science, Division of Materials Science and Engineering at Stanford under contract No. DE-AC02-76SF00515. EP was partially supported by the Koret Foundation.References
@@
- [3] O.J. Clark, M.J. Neat, K. Okawa, L. Bawden, I. Marković, F. Mazzola, J. Feng, V. Sunko, J.M. Riley, W. Meevasana, J. Fujii, I. Vobornik, T.K. Kim, M. Hoesch, T. Sasagawa, P. Wahl, M.S. Bahramy, and P.D.C. King,“Fermiology and Superconductivity of Topological Surface States in PdTe,” Phys. Rev. Lett. 120, 156401 (2018).
@@
- [6] Sebastian Schimmel, Yanina Fasano, Sven Hoffmann, Joaquín Puig, Grigory Shipunov, Danny Baumann, Saicharan Aswartham, Bernd Büchner, and Christian Hess, “High- surface superconductivity in topological Weyl semimetal t-PtBi,” arXiv:2302.08968 [cond-mat.supr-con] (2023).
@@
- [38] Xinyang Zhang, Mark Zic, Dong Chen, Chandra Shekhar, Claudia Felser, Ian R. Fisher, and Aharon Kapitulnik, “Vortex phase diagram of kagome superconductor CsVSb,” arXiv:2306.13297 [cond-mat.supr-con] (2023).
@@
- [44] Jing Wang, Xuliang Chen , Yonghui Zhou , Chao An, Ying Zhou, Chuanchuan Gu , Mingliang Tian, and Zhaorong Yang, “Pressure-induced superconductivity in trigonal layered PtBi with triply degenerate point fermions,” Phys. Rev. B 103, 014507 (2021).
@@
- [47] E. R. Schemm, W. J. Gannon, C. M. Wishne, W. P. Halperin, and A. Kapitulnik, “Observation of broken time-reversal symmetry in the heavy-fermion superconductor UPt,” Science 345, 190-193 (2014).
- [48] I. M. Hayes, D. S. Wei, T. Metz, J. Zhang, Y. S. Eo, S. Ran, S. R. Saha, J. Collini, N. P. Butch, D. F. Agterberg, A. Kapitulnik, and J. Paglione, “Multicomponent superconducting order parameter in UTe,” Science 373, 797-801 (2021).
@@
SUPPLEMENTAL INFORMATION
Possible Unconventional Surface Superconductivity in the Half-Heuslers LuPtBi and YPtBi
Eylon A. Persky, Alan Fang, Carolina Adamo, Eli Levenson-Falk, Qi I. Yang, Chandra Shekhar, Claudia Felser, Binghai Yan, and Aharon Kapitulnik
Geballe Laboratory for Advanced Materials, Stanford University, Stanford, California 94305, USA
Stanford Institute for Materials and Energy Sciences, SLAC National Accelerator Laboratory, 2575 Sand Hill Road, Menlo Park, California 94025, USA
Department of Applied Physics, Stanford University, Stanford, California 94305, USA
Department of Physics, Stanford University, Stanford, California 94305, USA
Max Planck Institute for Chemical Physics of Solids, 01187 Dresden, Germany.
Department of Condensed Matter Physics, Weizmann Institute of Science, Rehovot, Israel
S1 Crystal growth and sample fabrication
Single crystals of YPtBi and LuPtBi were grown by self-flux solution growth method, whereas Bi provides a flux medium. For example, the stoichiometry quantities of freshly polished piece of elements Lu, Pt and Bi of purity in the 0.3:0.3:10 atomic ratio were put in tantalum crucible and sealed in dry quartz ampoule under 3 mbar partial pressure of argon. The mixture containing ampoule was heated at a rate of 100 K/h up to 1473 K and followed by 12 hours soaking. For crystal growth, the temperature was slowly reduced by 2 K/h to 873 K and removed extra Bi flux by decanting the ampoule at 873 K. In this method, we get 2-3 mm regular triangle size of crystals, and their preferred growth orientation is (111), which was confirmed by Laue diffraction. The general method of crystals growth has been followed from literatures [29, 30]. The composition and structure were checked by energy dispersive X-ray analysis and Laue X-ray diffraction, respectively. The lattice parameters of the cubic structure are 6.574, which are consistent with previous reports [11, 14]. Energy dispersive x-ray spectrometry gives atomic percentages 33.6: 33.0: 33.4 3.0 for Lu: Pt: Bi confirming the stoichiometric ratio of the chemical compositions. Similar procedure and results were obtained for YPtBi.
S2 Surface Properties and Morphology
To understand the morphology and chemistry of the surface of the crystals we used X-ray Photoelectron Spectroscopy (XPS) that allows for high sensitivity elemental surface composition measurement and high resolution binding energy chemical shift measurement of solid samples under high vacuum. With typical measurement depth of about a nanometer, pristine crystals show traces of carbon, oxygen, and bismuth on the surface. However, after a several seconds of ion milling (equivalent to a few nm), all carbon traces are gone and all three elements: Lu, Pt and Bi seem to appear with the correct 1:1:1 ratio, together with traces of oxygen. This results does not changwee after subsequent milling of several more nanometers of the surface, confirming that no other elements exist in the samples.
S3 Surface cleaning procedure using STM tip
The ease by which the fresh surface oxidizes suggests that ex-situ milling and then loading the sample into the STM may not solve the surface contamination problem. To overcome this issue, we performed in-situ surface cleaning by using the STM itself to remove the top few layers and expose deeper layers, presumably reaching the true surface layer of the crystal. In a typical conditions of this process, performed in-situ at helium temperature under UHV conditions, the STM tip junction resistance was set by a bias of 20-50 mV and current of 100 pA. Then voltage pulses of a few volts were applied until zero bias conductance was minimized. Occasionally, we found regions of area of nm, which were flat and untilted to within a few angstroms, yet had an indistinct but reproducible topography (no atomic corrugations above 10’s of pm were seen, most likely due to tip bluntness). Since we focused on the temperature range below , the quality of spectra were gauged by how close the DOS reached zero at zero bias, as expected for gapped spectra. A similar method was previously used to expose superconducting regions in vacuum-cleaved YPtBi crystals by Baek et al. [32], although no particular effort was made to study the topological (111) direction of the crystals. Notably our (111) starting surface, while showing “lumpy” topography similar to [32], was overall much flatter, with only 0.3 nm height variations over 50 nm regions (as opposed to .1 nm in [32]).
S4 STM Measurements
As noted in the main text, we performed measurements on a modified UNISOKU Co., USM1300SHeTM system with Ultra High Vacuum (UHV), 350 mK base temperature, and magnetic field capabilities.
S5 Kerr Effect Measurements
High-resolution measurements of Kerr rotation were performed using a fiber-based zero-area loop Sagnac interferometer as first described in Ref. [33]. Following the main components shown in Fig. S4, two linearly polarized counterpropagating beams comprising the interferometer are isolated along the fast and slow axes of 10 m of polarization-maintaining fiber and are modulated in phase with a frequency that matches the transit time of light through the interferometer. The linearly polarized light traveling along each of these paths passes through a quarter-waveplate and is converted into circularly polarized light just above the sample. In the absence of magnetic field, the apparatus is completely reciprocal by symmetry except for the sample. Thus, upon reflection, one branch of the interferometer acquires a phase shift of , while its orthogonally polarized counterpart acquires an opposite phase shift of . The two phase shifts are added at the detector and extracted from the ratio of the first and second harmonics of the modulated signal.
S6 Ab-initio calculations of the surface band structure of LuPtBi
Ab initio calculations were performed within the density-funcitonal theory framework with the generalized gradient approximation, which is implemented into the VASP package [54]. A 54-atomic-layer-thick slab model was employed to simulate the LuPtBisurface, in which the top surface is terminated by Bi and the bottom surface is by Lu. A vacuum layer with more than 50 Åthick is employed in the -direction to avoid inter-slab interactions. A grid of 1212 is used for the 2D Brillouin zone integration. Spin-orbit coupling was included in all calculations.
In particular, the Bi-termination surface is shown in Fig. S5, exhibiting a van-Hove singularity (VHS) at eV near the M-point, about 5% offset along the line from M to . 6 such VHS points appear in the first Brillouin zone.

### v2

@@ Acknowledgements.
We thank Ronny Thomale and Jörg Schmalian for illuminating discussions. This work was supported by the U. S. Department of Energy (DOE) Office of Basic Energy Science, Division of Materials Science and Engineering at Stanford under contract No. DE-AC02-76SF00515 and by the National Science Foundation Award Number: 2307132. EP was partially supported by the Koret Foundation.References
@@
- [3] O.J. Clark, M.J. Neat, K. Okawa, L. Bawden, I. Marković, F. Mazzola, J. Feng, V. Sunko, J.M. Riley, W. Meevasana, J. Fujii, I. Vobornik, T.K. Kim, M. Hoesch, T. Sasagawa, P. Wahl, M.S. Bahramy, and P.D.C. King,“Fermiology and Superconductivity of Topological Surface States in PdTe2,” Phys. Rev. Lett. 120, 156401 (2018).
@@
- [6] Sebastian Schimmel, Yanina Fasano, Sven Hoffmann, Joaquín Puig, Grigory Shipunov, Danny Baumann, Saicharan Aswartham, Bernd Büchner, and Christian Hess, “High- surface superconductivity in topological Weyl semimetal t-PtBi2,” arXiv:2302.08968 [cond-mat.supr-con] (2023).
@@
- [29] Y.Z. Zhou, J. Chen, Z.X. Li, J. Luo, J. Yang, Y.F. Guo, W.H. Wang, R. Zhou, and Guo-ging Zheng, “Antiferromagnetic Spin Fluctuations and Unconventional Superconductivity in Topological Superconductor Candidate YPtBi Revealed by 195Pt-NMR,” Phys. Rev. Lett. 130, 266002 (2023).
- [30] J. C. Souza, M. V. Ale Crivillero, H. Dawczak-Debicki, Andrzej Ptok, P. G. Pagliuso, and S. Wirth, “Tuning the topological character of half-Heusler systems: A comparative study on YTBi (T=Pd,Pt),” Phys. Rev. B 108, 165154 (2023).
@@
- [40] Xinyang Zhang, Mark Zic, Dong Chen, Chandra Shekhar, Claudia Felser, Ian R. Fisher, and Aharon Kapitulnik, “Vortex phase diagram of kagome superconductor CsV3Sb5,” arXiv:2306.13297 [cond-mat.supr-con] (2023).
@@
- [46] Jing Wang, Xuliang Chen , Yonghui Zhou , Chao An, Ying Zhou, Chuanchuan Gu , Mingliang Tian, and Zhaorong Yang, “Pressure-induced superconductivity in trigonal layered PtBi2 with triply degenerate point fermions,” Phys. Rev. B 103, 014507 (2021).
@@
- [49] E. R. Schemm, W. J. Gannon, C. M. Wishne, W. P. Halperin, and A. Kapitulnik, “Observation of broken time-reversal symmetry in the heavy-fermion superconductor UPt3,” Science 345, 190-193 (2014).
- [50] I. M. Hayes, D. S. Wei, T. Metz, J. Zhang, Y. S. Eo, S. Ran, S. R. Saha, J. Collini, N. P. Butch, D. F. Agterberg, A. Kapitulnik, and J. Paglione, “Multicomponent superconducting order parameter in UTe2,” Science 373, 797-801 (2021).
@@
- [52] Troels A. Bojesen, Egor Babaev, and Asle Sudbø, ‘Time reversal symmetry breakdown in normal and superconducting states in frustrated three-band systems,” Phys. Rev. B 88, 220511(R) (2013).
- [53] Troels A. Bojesen, Egor Babaev, and Asle Sudbø, ‘Phase transitions and anomalous normal state in superconductors with broken time-reversal symmetry,” Phys. Rev. B 89, 104509 (2014).
@@
- [55] P. T. How and S. K. Yip, “Absence of Ginzburg-Landau mechanism for vestigial order in the normal phase above a two-component superconductor,” Phys. Rev. B 107, 104514 (2023).
@@
- [58] Jason K. Kawasaki, Abhishek Sharan, Linda I. M. Johansson, Martin Hjort, Rainer Timm, Balasubramanian Thiagarajan, Brian D. Schultz, Anders Mikkelsen, Anderson Janotti, and Chris J. Palmstrøm, “A simple electron counting model for half-Heusler surfaces,” Sci. Adv.4,eaar5832(2018).